# Feigenbaum universality in subcritical Taylor-Couette flow

### v1

@@ 2 Formulation of the problem
We consider an incompressible Newtonian fluid of kinematic viscosity $\nu$, filling the gap between two infinitely long concentric rotating cylinders. The angular velocities of the inner and outer cylinders, of radii $r^{*}_{\rm{i}}$ and $r^{*}_{\rm{o}}$, are denoted as $\Omega_{\rm{i}}$ and $\Omega_{\rm{o}}$, respectively. A complete set of independent dimensionless parameters characterising the problem are the radius ratio $\eta=r^{*}_{\rm{i}}/r^{*}_{\rm{o}}$ and the two Reynolds numbers $R_{i}=dr^{*}_{\rm{i}}\Omega_{\rm{i}}/\nu$ and $R_{o}=dr^{*}_{\rm{o}}\Omega_{\rm{o}}/\nu$, where $d=r^{*}_{\rm{o}}-r^{*}_{\rm{i}}$ is the gap.
@@
All computations have been run in this domain. The same resolution of $[0,50]\times[-8,8]\times[-8,8]$ modes as used in W22 has been employed throughout.
@@
The system possesses translational invariance in $\xi$ and $\zeta$. We define the entire set of possible spectral coefficients as the phase space. All coefficient vectors that belong to the group orbit induced by translation of a same velocity field represent the same solution. We have systematically factored out the group orbit invariance employing the method of slices (Budanur et al., 2015), using the same template as Wang et al. (2023). To analyse the period doubling cascade within the framework of discrete-time dynamical systems, we have devised a Poincaré section
@@
in phase space $\mathbb{X}$, which consists of all possible sliced spectral expansion coefficients vector $\mathbf{a}$. The condition defining $\Sigma$ is based on the equality of inner and outer cylinder torque (first condition) and the sign of the rate of change of their difference (second condition) to discard reverse crossings. It is easy to check from the governing equations that for statistically steady states the time average of $\tau_{i}$ must coincide with that of $\tau_{o}$. Consequently, long-lived (including permanent) time-dependent solutions are bound to regularly fulfil the condition $\tau_{\rm{i}}=\tau_{\rm{o}}$, a property that comes in handy in defining a robust Poincaré section. To simplify notation, we will hereafter call $\tau=\tau_{i}=\tau_{o}$ the value of the torque, inner or outer indistinctly, on $\Sigma$.
@@ 3 The period doubling cascade
Emulating the influential experiments by Andereck et al. (1986), we employ a radius ratio $\eta=0.883$ ($r\in[r_{i},r_{o}]=[7.547,8.547]$). The outer cylinder Reynolds number is fixed at $R_{o}=-1200$. For sufficiently large $R_{i}$, the centrifugal instability of the base flow develops into full-fledged turbulence. A reduction of $R_{i}$ while still in the supercritical region has the flow re-organise into a pattern of alternated laminar and turbulent winding helical bands, commonly known as the spiral turbulence regime (Coles, 1965; Andereck et al., 1986; Meseguer et al., 2009b; Dong, 2009). At this particular value of $R_{o}$, the CCF is linearly stable for $R_{i}\lesssim 447.35$, but a number of subcritical nonlinear equilibrium states persists below this threshold (Meseguer et al., 2009a; Deguchi et al., 2014, and W22), which are believed to act as precursors of spiral turbulence. Drawing from the helix angle of the turbulent stripes observed experimentally, W22 employed the domain and coordinate system shown in figure 1a to seek for minimal flow unit solutions that are compatible with the tilt of the banded pattern. Subharmonic instabilities of some such solutions were proposed as the mechanism that might be responsible for the spatial intermittency characterising spiral turbulence. Within this minimal parallelogram-annular periodic domain, which is capable of sustaining turbulence at sufficiently high $R_{i}$, W22 identified a stable finite amplitude travelling wave (shown in figure 1b) that coexists with the stable CCF. This drifting rotating wave, DRW for short, embodies all the essential elements of the self-sustaining mechanism (Wang et al., 2007; Hall & Sherwin, 2010) and plays a central role in the transition process. In W22, two periodic solutions, P1, arising from a Hopf bifurcation of DRW, and P2, originated at a period-doubling bifurcation of P1, were also identified as the first steps in the route to chaos. This sequence of bifurcations suggested a period-doubling cascade as the most plausible scenario for the onset of chaotic dynamics, but the issue was not pursued further.
@@ 3.1 The onset of the period doubling cascade
We now look into the bifurcation sequence the system undergoes in the range $R_{i}\in[395.43,395.79]$ eventually leading to sustained chaotic dynamics. Since we will only be varying $R_{i}$, we shall drop the subindex and simply write $R$ to ease notation.
@@
DRW, a relative equilibrium, is characterised by a solid-body motion composing axial translation and azimuthal rotation with constant phase velocities. As a result, the space-time diagram exhibits the repetition of a periodic pattern. The same line probe produces a quasi-periodic space-time diagram for P1, as shown in figure 3bi. P1 is a relative periodic orbit and, therefore, requires suitable shifts in both the $z$ and $\theta$ directions to align the flow fields after every one period to reveal the space-time invariance. The effects of the drift can be dismissed by attaching the line probe to a moving reference frame defined with the method of slices (Budanur et al., 2015; Wang et al., 2023). In this reference frame, the space-time diagrams of DRW and P1 (figures 3aii and 3bii) are greatly simplified, the former appearing as time-independent and the latter as purely time-periodic.
@@
DRW emerges from a saddle-node bifurcation (SN1) at $R\simeq 391.50$ and, leaving aside subharmonic instabilities, the nodal (upper) branch (the one represented in figure 1b and the square in figures 2a,c) remains stable to perturbations fitting the domain until the advent of a supercritical Hopf bifurcation (H) at $R\simeq 392.85$. As reported by W22, this is the only known case of a subcritically stable non-trivial solution in Taylor-Couette flow (note that here we are only interested in the stability to perturbations that fit within the periodic box). Stable/unstable solutions are denoted by solid/dashed lines. The PNK method must, of course, be used to compute unstable solution branches. A stable relative periodic orbit (P1, blue line) bifurcates from DRW at H. All time-dependent solutions will be represented, from now on, through the collection of intersection points on the Poincaré section. The P1 branch becomes unstable in a period doubling bifurcation (PD1), whence a branch of stable period-doubled relative periodic orbits (P2, pair of green lines) emerges. The P2 loses stability, in turn, in a second period-doubling bifurcation (PD2) issuing a branch of period-4 solutions. The $R$ value for which figures 2 and 3 were computed corresponds to just short of PD2, hence the instability of DRW and P1, in contrast with P2, which is stable.
The remainder of the period-doubling cascade and the onset of chaotic dynamics at larger values of $R$ is depicted in figure 4b. Only stable solutions are shown, so DNS is all that has been required to produce the diagram. Estimating $\delta_{F}$ from the approximate location on the bifurcation diagram of the period doubling points may seem a straightforward task. However, confirmation of Feigenbaum universality by brute force poses challenges, even for simple systems as the logistic map. To begin with, the parameter spacing between consecutive period-doubling bifurcations shrinks very fast as one progresses along the cascade, demanding the computation of bifurcation points with an ever increasing number of significant digits. Moreover, the orbits in the immediate vicinity of a period doubling point, which are required for the accurate estimation of the bifurcations, are close to neutrally stable and, therefore, take massive computational time to reach convergence with DNS. Unfortunately, employing the PNK method becomes impractical for solutions of increasingly long periods. Both complications combined render the careful estimation of Feigenbaum’s first constant extremely hard. A rigorous and systematic analysis as detailed in the coming section becomes thus necessary.
@@ 3.2 Determination of period doubling points
To analyse the period-doubling cascade, it is most convenient to focus on the sequence of ${\bf a}$ on $\Sigma$. In the framework of dynamical systems theory, this corresponds to investigating the properties of the Poincaré map, also known as first return map. As we shall explain later, it is not necessary to monitor the full set of coefficients, and the sequence of torque values provides (nearly) all necessary information in the long run, once initial transients have been overcome. Let $\tau(\ell)$, $\ell=1,2,3,\dots$ be the corresponding sequence of torque values on $\Sigma$, with $\ell$ an index recording the chronological order. As an example, figure 5a shows $\tau(\ell)$ for the P4 solution (red circles) at $R=395.711$ (between PD2 and PD3) and for the P8 solution (blue) at $R=395.719$ (between PD3 and PD4).
@@
Both solutions are stable and well converged with DNS, such that the respective series sequentially repeat, always in the same order, the same four (P4) and eight (P8) distinct values, as indicated by the numbered horizontal dashed lines. The effect of the period-doubling bifurcation PD3 can thus be portrayed as doubling point $j$ of the P4 discrete-time orbit into points $j$ and $j+4$ of the P8 cycle, the latter two points having sprung from the former and drifted away in opposite directions. The unstable P1 (cyan) and P2 (green) orbits in the immediate vicinity of PD3 are shown for reference, to help generalise the rule that relates points $j$ and $j+N/2$ of the period-$N$ orbit to point $j$ of the period-$N/2$ orbit from which they bifurcate. A phase map analogous to that of figure 2c is shown in figure 5b, but with the region where $\Sigma$ is pierced by the orbits duly magnified to further clarify where the P4 and P8 cycles stand in relation to P1 and P2.
The $n$th period doubling bifurcation PDn can be readily analysed by splitting the $\tau(\ell)$ sequences of interest in $J=2^{n}$ separate subsequences
@@
To nail down PD3, a collection of DNS runs traversing the bifurcation point, i.e., between $R=395.711$ and $395.719$, are required. The amplitude of P8 drops fast as the parameter $R$ is decreased towards PD3, such that discerning it from P4 becomes increasingly difficult. In addition, the dynamics become despairingly slow as the bifurcation point is approached from either side, rendering convergence with DNS downright impracticable. Figures 6a and 6b depict the torque subsequences at $R=395.7116$ and $R=395.7122$, respectively, below and above PD3 but moderately close to the bifurcation point.
@@
The eight $\lambda_{j}^{8}$ multipliers, shown in figure 7b, behave as expected across a period doubling bifurcation. The eight individual fits (15) to the eight individual $\tau_{j}^{8}$ subsequences produce nearly identical values of $\lambda_{j}^{8}$ at any given $R$, hence the multicoloured filled circles, resulting from the superposition of the eight oriented sectored disks. The unique estimate of the multiplier at any given $R$, obtained as the arithmetic mean $\lambda_{8}=\langle\lambda_{j}^{8}\rangle$, increases towards unity as the period doubling point is approached from either side. Since the solution above PD3 is indeed a P8, $\lambda_{8}$ is directly its multiplier. This is not so for the P4 solution below PD3. Plots analogous to those in figure 6a but sampling with $J=4$ show that convergence onto the four asymptotic branches follows an oscillatory pattern. The multiplier of the P4 orbit in the vicinity of PD3 must be either estimated from fits to the torque subsequences sampled with $J=4$ or indirectly from $\lambda_{8}$ according to $\lambda_{4}=\langle\lambda_{j}^{4}\rangle=-\sqrt{\lambda_{8}}$. Either way, the multiplier of the stable P4 in the vicinity of PD3 is real, negative, and on the way of leaving the unit circle of the complex plane through $\lambda=-1$ as $R_{3}$ is approached from below (black disks in figure 7b), consistent with the period doubling nature of the bifurcation.
@@ 4 Verification of Feigenbaum universality
Successive magnifications, presented in 8b and 8c, exhibit branch structures very similar to that of the full cascade of 8a. Yet, at this stage, there is no guarantee that the universality observed here is of the same kind as discovered by Feigenbaum.
In order to ascertain Feigenbaum universality, we have estimated the sequence of critical $R_{n}$ values along the period doubling cascade up to order $n=7$ to very high accuracy, using the method described in §3.2. The results are summarised in table 2.
@@
Table 2: Confirmation of Feigenbaum universality along the period doubling cascade. The values of the critical inner cylinder Reynolds number $R_{n}$ at the $n$th period doubling bifurcation point PDn are used to compute the $n$th estimate to Feigenbaum’s first constant, $\delta_{n}$, according to (17). An estimate to the second constant, $\alpha_{n}$, is computed from central branch torque values following (24). Lower bounds to the uncertainties in the $R_{n}$, $\delta_{n}$, $\alpha_{n}$ and $R_{\infty}$ parameters have been rigorously estimated from the covariance matrices of the various fits involved in the process, combined with standard error propagation theory (note that we have not taken into account the error of the fit (15), as conducting a systematic study is difficult). The last row of the table corresponds to the accumulation point $R_{\infty}$ estimated by (23), and the actual values of Feigenbaum constants, $\delta_{\rm F}$ and $\alpha_{\rm F}$.
@@
which provide the best estimate of Feigenbaum’s first constant $\delta_{\rm F}$ that may be obtained from data up to the $n$th period doubling. The trend of $\delta_{n}$ exhibits evident signs of conforming to Feigenbaum universality, that is, $\delta_{n}\rightarrow\delta_{\rm F}$ as $n\rightarrow\infty$. It is instructive to compare these results with those in table 1. Limiting the analysis to $n=4$, as customarily done in the fluid dynamics literature, does not guarantee convergence to even the second significant digit. Securing three digits precision is an arduous task that typically requires the accurate determination of at least $n=7$ period doubling bifurcation points.
The $R_{n}$ sequence is expected to converge onto the accumulation point, $R=R_{\infty}$. This point may be estimated by assuming that all period doubling bifurcations of order higher than available are perfectly Feigenbaum-universal and, therefore, satisfy (17) with the left hand side replaced by $\delta_{\rm F}$. However, this naive approach leads to systematic error. To obtain a better estimate of $R_{\infty}$ and, at the same time, bound the error, we first note that
@@
The constant $\tilde{\delta}_{n}$ is, of course, unknown and may only be estimated. For $n$ sufficiently large, the sequence $\delta_{\ell}$ for all $\ell>n$ is expected to approach $\delta_{\rm F}$ monotonically from above. Under this assumption, inspection of (21) provides the inequalities $\delta_{\rm F}<\delta_{n+1}<\tilde{\delta}_{n}<\delta_{n}$. Accordingly, upper and lower bounds for $R_{\infty}$ can be obtained as
@@
If the manifold has dimension one, the coefficients $\mathbf{a}$ on the Poincaré section $\Sigma$ introduced in (11) can be uniquely parameterised by the torque, such that the first return map $f$ defined by $\tau(\ell+1)=f(\tau(\ell))$ contains all the long-term properties of the Poincaré map. The aperiodicity of the orbit becomes apparent when plotting $\tau(\ell+1)$ as a function of $\tau(\ell)$ (green dots in figure 10b). The structure of the map on $\Sigma$ is, however, not clear, as the $\Sigma$-crossings of the orbit are sparsely distributed (see figure 10a). Perturbing the orbit, dropping the first initial transients, and plotting the remainder of the transients, conveniently assists in populating the map more densely (black points). All the points thus obtained fall within a narrow, nearly one-dimensional band, seemingly connected by a smooth curve. This indicates that the discrete-time dynamics on $\Sigma$ in the vicinity of the accumulation point can be well approximated by a one-dimensional map $f$. The function $f$ is unimodal, and its smooth extremum (the minimum) admits a quadratic approximation. Note that unimodal maps with non-quadratic extrema provide universality constants that deviate from Feigenbaum’s well-known values (Briggs, 1991).
For period-doubling cascades in one-dimensional unimodal maps, the second Feigenbaum constant, $\alpha_{\rm F}$, is often estimated from the central branch of the bifurcation diagram, which is generated by the periodic point that is nearest to the extremum of the map. Since producing accurate bifurcation diagrams of cascades undergone by fluid flow systems is an arduous task, Feigenbaum (1979) used instead the frequency spectrum, recorded at a fixed value of the parameter, of a natural convection experiment. In the subsequent studies summarised in table 1, the estimation of $\alpha_{\rm F}$ has not been conducted.
On the other hand, we have a highly accurate bifurcation diagram available, making the direct estimation of the second constant possible. In the torque sequences presented so far, we have indexed the $N$ distinct points visited by every periodic solution PN as $j=1,2,\dots N$ such that $j=1$ is the central branch point (the black dots in figure 9). Thus, using the torque values of the central branch at PDm, $\hat{\tau}_{m}=\tau_{1}^{2^{m}}(R_{m})$, we can compute the ratios
@@
to provide estimates for the $\alpha_{\rm F}$ appearing in the Feigenbaum-Cvitanović functional equation. The values of $\alpha_{n}$ thus obtained from the bifurcation diagram, listed in table 2 alongside $\delta_{n}$, evince gradual convergence toward Feigenbaum universality.

### v2

@@ 2 Formulation of the problem
We consider an incompressible Newtonian fluid of kinematic viscosity $\nu$, filling the gap between two infinitely long concentric rotating cylinders (figure 1a). The angular velocities of the inner and outer cylinders, of radii $r^{*}_{\rm{i}}$ and $r^{*}_{\rm{o}}$, are denoted as $\Omega_{\rm{i}}$ and $\Omega_{\rm{o}}$, respectively. A complete set of independent dimensionless physical parameters characterising the problem are the radius ratio $\eta=r^{*}_{\rm{i}}/r^{*}_{\rm{o}}$ and the two Reynolds numbers $R_{i}=dr^{*}_{\rm{i}}\Omega_{\rm{i}}/\nu$ and $R_{o}=dr^{*}_{\rm{o}}\Omega_{\rm{o}}/\nu$, where $d=r^{*}_{\rm{o}}-r^{*}_{\rm{i}}$ is the gap.
@@
The domain was specifically designed for the earliest onset of nontrivial solutions while keeping compatibility with the large-scale tilt of spiral turbulence. For this reason, the bifurcation scenario we investigate is fully compatible with the overall structure of intermittent patterns experimentally observed in the counter-rotating regime of the Taylor-Couette system and, at the same time, can be reasonably expected to precede any other mechanism potentially leading to the onset of chaos. All computations have been run in this domain. The same resolution of $[0,50]\times[-8,8]\times[-8,8]$ modes as used in W22 has been employed throughout. The system of ordinary differential equations that results from spatial discretisation has dimension $O(10^{5})$.
@@
The system possesses translational invariance in $\xi$ and $\zeta$. We define the entire set of possible spectral coefficients as the phase space. All coefficient vectors that belong to the group orbit induced by translation of a velocity field represent the same solution. We have systematically factored out the group orbit invariance employing the method of slices (Budanur et al., 2015), using the same template as Wang et al. (2023). To analyse the period doubling cascade within the framework of discrete-time dynamical systems, we have devised a Poincaré section
@@
in phase space $\mathbb{X}$, which consists of all possible sliced spectral expansion coefficients vector $\mathbf{a}$. The condition defining $\Sigma$ is based on the equality of inner and outer cylinder torque (first condition) and the sign of the rate of change of their difference (second condition) to discard reverse crossings. It is easy to check from the governing equations that for statistically steady states the time average of $\tau_{i}$ must coincide with that of $\tau_{o}$. Consequently, long-lived (including permanent) time-dependent solutions are bound to regularly fulfil the condition $\tau_{\rm{i}}=\tau_{\rm{o}}$, a property that comes in handy in defining a robust Poincaré section. To simplify notation, we will hereafter call $\tau=\tau_{i}=\tau_{o}$ the value of the torque, whether inner or outer, on $\Sigma$.
There are other ways of sampling a continuous system to produce a discrete system. For example. Kreilos & Eckhardt (2012) and Gao et al. (2015) investigated the relationship between consecutive local maxima of the time series of some physical quantity. However, it is not easy to discern which minima/maxima are actually related to the underlying (pseudo-)periodicity of the solutions and which are simply incidental to the dynamics of the particular time signal used. Thus the use of the above tailor-made Poincaré section provides a more reliable approach.
@@ 3 The period doubling cascade
Emulating the influential experiments by Andereck et al. (1986), we employ a radius ratio $\eta=0.883$ ($r\in[r_{i},r_{o}]=[7.547,8.547]$). The outer cylinder Reynolds number is fixed at $R_{o}=-1200$. For sufficiently large $R_{i}$, the centrifugal instability of the base flow develops into full-fledged turbulence. A reduction of $R_{i}$ while still in the supercritical region has the flow re-organise into a pattern of alternating laminar and turbulent winding helical bands, commonly known as the spiral turbulence regime (Coles, 1965; Andereck et al., 1986; Meseguer et al., 2009b; Dong, 2009). At this particular value of $R_{o}$, the CCF is linearly stable for $R_{i}\lesssim 447.35$, but a number of subcritical nonlinear equilibrium states persists below this threshold (Meseguer et al., 2009a; Deguchi et al., 2014, and W22), which are believed to act as precursors of spiral turbulence. Drawing from the helix angle of the turbulent stripes observed experimentally, W22 employed the domain and coordinate system shown in figure 2a to seek minimal flow unit solutions that are compatible with the tilt of the banded pattern. Subharmonic instabilities of some such solutions were proposed as the mechanism that might be responsible for the spatial intermittency characterising spiral turbulence. Within this minimal parallelogram-annular periodic domain, which is capable of sustaining turbulence at sufficiently high $R_{i}$, W22 identified a stable finite amplitude travelling wave (shown in figure 2b) that coexists with the stable CCF. This drifting rotating wave, DRW for short, embodies all the essential elements of the self-sustaining mechanism (Wang et al., 2007; Hall & Sherwin, 2010) and plays a central role in the transition process. In W22, two periodic solutions, P1, arising from a Hopf bifurcation of DRW, and P2, originating at a period-doubling bifurcation of P1, were also identified as the first steps in the route to chaos. This sequence of bifurcations suggested a period-doubling cascade as the most plausible scenario for the onset of chaotic dynamics, but the issue was not pursued further.
@@ 3.1 The onset of the period doubling cascade
We now look into the bifurcation sequence the system undergoes in the range $R_{i}\in[395.43,395.79]$ eventually leading to sustained chaotic dynamics. Since we will only be varying $R_{i}$, we shall drop the subscript and simply write $R$ to ease notation.
@@
DRW, a relative equilibrium, is characterised by a solid-body motion composing axial translation and azimuthal rotation with constant phase velocities. As a result, the space-time diagram exhibits the repetition of a periodic pattern. The same line probe produces a quasi-periodic space-time diagram for P1, as shown in figure 4bi. P1 is a relative periodic orbit and, therefore, requires suitable shifts in both the $z$ and $\theta$ directions to align the flow fields after every one period to reveal the space-time invariance. The effects of the drift can be suppressed by attaching the line probe to a moving reference frame defined with the method of slices (Budanur et al., 2015; Wang et al., 2023). In this reference frame, the space-time diagrams of DRW and P1 (figures 4aii and 4bii) are greatly simplified, the former appearing as time-independent and the latter as purely time-periodic.
@@
DRW emerges from a saddle-node bifurcation (SN1) at $R\simeq 391.50$ and, leaving aside subharmonic instabilities, the nodal (upper) branch (the one represented in figure 2b and the square in figures 3a,c) remains stable to perturbations fitting the domain until the advent of a supercritical Hopf bifurcation (H) at $R\simeq 392.85$. As reported by W22, this is the only known case of a stable non-trivial solution in the subcritical parameter regime of Taylor-Couette flow. Note that here we are only interested in the stability to perturbations that fit within the periodic box. Stable/unstable solutions are denoted by solid/dashed lines. The PNK method must be used to compute unstable solution branches. A stable relative periodic orbit (P1, blue line) bifurcates from DRW at H. All time-dependent solutions will be represented, from now on, through the collection of intersection points on the Poincaré section. The P1 branch becomes unstable in a period doubling bifurcation (PD1), whence a branch of stable period-doubled relative periodic orbits (P2, pair of green lines) emerges. The P2 loses stability, in turn, in a second period-doubling bifurcation (PD2) issuing a branch of period-4 solutions. The $R$ value for which figures 3 and 4 were computed corresponds to just short of PD2, hence the instability of DRW and P1, in contrast with P2, which is stable.
The remainder of the period-doubling cascade and the onset of chaotic dynamics at larger values of $R$ is depicted in figure 5b. Only stable solutions are shown, so DNS is all that has been required to produce the solution branches in the diagram. Checking agreement with $\delta_{F}$ from the approximate location on the bifurcation diagram of the period doubling points may seem a straightforward task. However, confirmation of Feigenbaum universality by brute force poses challenges, even for simple systems such as the logistic map. To begin with, the parameter spacing between consecutive period-doubling bifurcations shrinks very fast as one progresses along the cascade, demanding the computation of bifurcation points with an ever increasing number of significant digits. Moreover, the orbits in the immediate vicinity of a period doubling point, which are required for the accurate estimation of the bifurcations, are close to neutrally stable and, therefore, take massive computational time to reach convergence with DNS. Unfortunately, employing the PNK method becomes impractical for solutions of increasingly long periods. The inability of the method to discriminate between stable and unstable orbits, combined with the fact that stable orbits in a period doubling cascade coexist with all unstable orbits of lower period, which are favoured unless very close initial guesses are produced in advance, curbs any advantage one might have expected to achieve from using PNK. Both complications combined render the careful appraisal of agreement with Feigenbaum’s first constant extremely hard. A rigorous and systematic analysis as detailed in the coming section becomes thus necessary.
@@ 3.2 Determination of period doubling points
To analyse the period-doubling cascade, it is most convenient to focus on the sequence of ${\bf a}$ on $\Sigma$. In the framework of dynamical systems theory, this corresponds to investigating the properties of the Poincaré map, also known as first return map. As we shall explain later, it is not necessary to monitor the full set of coefficients, and the sequence of torque values provides (nearly) all necessary information in the long run, once the initial transients are over. Let $\tau(\ell)$, $\ell=1,2,3,\dots$ be the corresponding sequence of torque values on $\Sigma$, with $\ell$ an index recording the chronological order. As an example, figure 6a shows $\tau(\ell)$ for the P4 solution (red circles) at $R=395.711$ (between PD2 and PD3) and for the P8 solution (blue) at $R=395.719$ (between PD3 and PD4).
@@
Both solutions are stable and well converged with DNS, such that the respective series sequentially repeat, always in the same order, the same four (P4) and eight (P8) distinct values, as indicated by the numbered horizontal dashed lines. The effect of the period-doubling bifurcation PD3 can thus be portrayed as doubling point $j$ of the P4 discrete-time orbit into points $j$ and $j+4$ of the P8 cycle, the latter two points having sprung from the former and drifted away in opposite directions. The unstable P1 (cyan) and P2 (green) orbits in the immediate vicinity of PD3 are shown for reference, to help generalise the rule that relates points $j$ and $j+N/2$ of the period-$N$ orbit to point $j$ of the period-$N/2$ orbit from which they bifurcate. A phase map analogous to that of figure 3c is shown in figure 6b, but with the region where $\Sigma$ is pierced by the orbits suitably magnified to further clarify where the P4 and P8 cycles stand in relation to P1 and P2.
The $n$th period doubling bifurcation PDn can be readily analysed by splitting the $\tau(\ell)$ sequences of interest into $J=2^{n}$ separate subsequences
@@
To nail down PD3, a collection of DNS runs traversing the bifurcation point, i.e., between $R=395.711$ and $395.719$, are required. The amplitude of P8 drops fast as the parameter $R$ is decreased towards PD3, such that distinguishing it from P4 becomes increasingly difficult. In addition, the dynamics become despairingly slow as the bifurcation point is approached from either side, rendering convergence with DNS downright impracticable. Figures 7a and 7b depict the torque subsequences at $R=395.7116$ and $R=395.7122$, respectively, below and above PD3 but moderately close to the bifurcation point.
@@
The eight $\lambda_{j}^{8}$ multipliers, shown in figure 8b, behave as expected across a period doubling bifurcation. The eight individual fits (15) to the eight individual $\tau_{j}^{8}$ subsequences produce nearly identical values of $\lambda_{j}^{8}$ at any given $R$, hence the multicoloured filled circles, resulting from the superposition of the eight oriented sectored disks. The unique estimate of the multiplier at any given $R$, obtained as the arithmetic mean $\lambda_{8}=\langle\lambda_{j}^{8}\rangle$, increases towards unity as the period doubling point is approached from either side. Since the solution above PD3 is indeed a P8, $\lambda_{8}$ is directly its multiplier. This is not so for the P4 solution below PD3. Plots analogous to those in figure 7a but sampling with $J=4$ show that convergence onto the four asymptotic branches follows an oscillatory pattern. The multiplier of the P4 orbit in the vicinity of PD3 must be either computed from fits to the torque subsequences sampled with $J=4$ or indirectly from $\lambda_{8}$ according to $\lambda_{4}=\langle\lambda_{j}^{4}\rangle=-\sqrt{\lambda_{8}}$. Either way, the multiplier of the stable P4 in the vicinity of PD3 is real, negative, and on the way of leaving the unit circle of the complex plane through $\lambda=-1$ as $R_{3}$ is approached from below (black disks in figure 8b), consistent with the period doubling nature of the bifurcation.
@@ 4 Verification of Feigenbaum universality
Successive magnifications, presented in 9b and 9c, exhibit branch structures very similar to that of the full cascade of 9a. Yet, at this stage, there is no guarantee that the universality observed here is of the same kind as discovered by Feigenbaum.
In order to ascertain Feigenbaum universality, we have computed the sequence of critical $R_{n}$ values along the period doubling cascade up to order $n=7$ to very high accuracy, using the method described in §3.2. The results are summarised in table 2.
@@
Table 2: Confirmation of Feigenbaum universality along the period doubling cascade. The values of the critical inner cylinder Reynolds number $R_{n}$ at the $n$th period doubling bifurcation point PDn are used to compute the $n$th approximation, $\delta_{n}$, to Feigenbaum’s first constant, according to (17). Approximations, $\alpha_{n}$, to the second constant, are computed from central branch torque values following (24). Lower bounds to the uncertainties in the $R_{n}$, $\delta_{n}$, $\alpha_{n}$ and $R_{\infty}$ parameters have been rigorously estimated from the covariance matrices of the various fits involved in the process, combined with standard error propagation theory (note that we have not taken into account the error of the fit (15), as conducting a systematic study is difficult). The last row of the table corresponds to the accumulation point $R_{\infty}$ estimated by (23), and the actual values of Feigenbaum constants, $\delta_{\rm F}$ and $\alpha_{\rm F}$.
@@
which provide the best approximation to Feigenbaum’s first constant $\delta_{\rm F}$ that may be obtained from data up to the $n$th period doubling. The trend of $\delta_{n}$ exhibits evident signs of conforming to Feigenbaum universality, that is, $\delta_{n}\rightarrow\delta_{\rm F}$ as $n\rightarrow\infty$. It is instructive to compare these results with those in table 1. Limiting the analysis to $n=4$, as customarily done in the fluid dynamics literature, does not guarantee convergence to even the second significant digit. Securing three digits precision is an arduous task that typically requires the accurate determination of at least $n=7$ period doubling bifurcation points.
The $R_{n}$ sequence is expected to converge onto the accumulation point, $R=R_{\infty}$. This point may be computed by assuming that all period doubling bifurcations of order higher than available are perfectly Feigenbaum-universal and, therefore, satisfy (17) with the left hand side replaced by $\delta_{\rm F}$. However, this naive approach leads to systematic error. To obtain a better estimate of $R_{\infty}$ and, at the same time, bound the error, we first note that
@@
The constant $\tilde{\delta}_{n}$ is unknown and may only be estimated. For $n$ sufficiently large, the sequence $\delta_{\ell}$ for all $\ell>n$ is expected to approach $\delta_{\rm F}$ monotonically from above. Under this assumption, inspection of (21) provides the inequalities $\delta_{\rm F}<\delta_{n+1}<\tilde{\delta}_{n}<\delta_{n}$. Accordingly, upper and lower bounds for $R_{\infty}$ can be obtained as
@@
If the manifold has dimension one, the coefficients $\mathbf{a}$ on the Poincaré section $\Sigma$ introduced in (11) can be uniquely parameterised by the torque, such that the first return map $f$ defined by $\tau(\ell+1)=f(\tau(\ell))$ contains all the long-term properties of the Poincaré map. The aperiodicity of the orbit becomes apparent when plotting $\tau(\ell+1)$ as a function of $\tau(\ell)$ (green dots in figure 11b). The structure of the map on $\Sigma$ is, however, not clear, as the $\Sigma$-crossings of the orbit are sparsely distributed (see figure 11a). Perturbing the orbit, dropping the first initial transients, and plotting the remainder of the transients, conveniently assists in populating the map more densely (black points). All the points thus obtained fall within a narrow, nearly one-dimensional band, seemingly connected by a smooth curve. This indicates that the discrete-time dynamics on $\Sigma$ in the vicinity of the accumulation point can be well approximated by a one-dimensional map $f$. Note that, in the range of $\tau$ for which $f$ appears to be multivalued (see the black points), the lower branch is never visited by the attractor and can, therefore, be ignored as regards its dynamics. Accordingly, the function $f$, with its leftmost branch duly trimmed, becomes unimodal, and its smooth extremum (the minimum) admits a quadratic approximation. Note that unimodal maps with non-quadratic extrema provide universality constants that deviate from Feigenbaum’s well-known values (Briggs, 1991).
For period-doubling cascades in one-dimensional unimodal maps, approximations of the second Feigenbaum constant, $\alpha_{\rm F}$, are often obtained from the central branch of the bifurcation diagram, which is generated by the periodic point that is nearest to the extremum of the map. Since producing accurate bifurcation diagrams of cascades undergone by fluid flow systems is an arduous task, Feigenbaum (1979) used instead the frequency spectrum, recorded at a fixed value of the parameter, of a natural convection experiment. In the subsequent studies summarised in table 1, no evaluation of agreement with $\alpha_{\rm F}$ was conducted.
On the other hand, we have a highly accurate bifurcation diagram available, rendering direct assessment of agreement with the second constant possible. In the torque sequences presented so far, we have indexed the $N$ distinct points visited by every periodic solution PN as $j=1,2,\dots N$ such that $j=1$ is the central branch point (the black dots in figure 10). Thus, using the torque values of the central branch at PDm, $\hat{\tau}_{m}=\tau_{1}^{2^{m}}(R_{m})$, we can compute the ratios
@@
to be compared against the $\alpha_{\rm F}$ appearing in the Feigenbaum-Cvitanović functional equation. The values of $\alpha_{n}$ thus obtained from the bifurcation diagram, listed in table 2 alongside $\delta_{n}$, evince gradual convergence toward Feigenbaum universality.